OpenStudy (firejay5):

Find the value of each variable. If your answer is NOT an integer, leave it in simplest radical form. Show work to get medal! :D

OpenStudy (firejay5):



OpenStudy (anonymous):

\[ 9 = 2x \]

OpenStudy (anonymous):

So \(x=9/2\)

OpenStudy (anonymous):

\[ y^2 = 9^2 + (9/2)^2 \]

OpenStudy (anonymous):

Can you find \(y\)?

OpenStudy (firejay5):

@wio Is this using Pythagorean Theorem?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Notice the \(^2\)s everywhere.

OpenStudy (firejay5):

Well I thought would ask